The speed of glycerin flowing in a 5.0 cm i.d. pipe is 0.54 m/s. Find the fluid‘s speed in a 3.0 cm i.d. pipe that connects with it, both pipes flowing full. Ans. 1.5 m/s

 


1
Expert's answer
2021-11-23T10:33:16-0500
"0.54(5)^2=(3)^2v\\v=1.5\frac{m}{s}"